Use the given degree of confidence and sample data to construct a confidence interval for the population mean u, assume that the population has a normal distribution. N=12,x=21.3, s=6.3, 99%
1
Expert's answer
2012-04-13T10:28:02-0400
Sample mean = x-bar = 21.3
ME = z*s/sqrt(n) = 2.5758*6.3/sqrt(12) = 4.6845
99% CI: 21.3-4.68 < u < 21.3+4.68, or (16.62,25.98).